OpenStudy (anonymous):

30 children r going on a trip it costs £5 including lunch Some go with their own lunch so they pay £3 The 30 children pay a total of £110 How many take their own lunch?

OpenStudy (anonymous):

20 take their own lunch. 20*3=60 5*10=50 60+50=110

OpenStudy (anonymous):

Thx

OpenStudy (anonymous):

trial and error

OpenStudy (anonymous):

20 take ther own

OpenStudy (anonymous):

I had it before but I forgot

OpenStudy (anonymous):

I didn't use trial and error. I guessed right the first time. It was just luck though.

OpenStudy (anonymous):

lucky it was trial and error when I did it though ;)